# On the Onset of Coherent Phonon Motion in Peierls-Distorted Antimony by Attosecond Transient Absorption

### v2

@@ Results
To examine the break-down of the two temperature model after the initial excitation by the pump pulse and the following carrier thermalization more closely, the measured transient absorption spectrum at early time delays are compared to the predicted spectrum from the TTM. As can be seen in Fig. 2a, the experimental data show discrepancies to the model due to the large number of non-thermalized carriers shortly after the pump pulse is over (12 fs). Compared to the model, a larger absorbance is found in a region 1 eV below the Fermi level, indicating that the pump excites carriers largely from the lower bands at the L- and Z-point (cf. Fig. 1a,b). On the other hand, at energies up to 0.5 eV below the Fermi level, there is less absorbance than expected from the model, indicating that fewer holes are initially in the highest valence bands. Above the Fermi level, a stronger negative change in absorbance is observed, peaking at 32.6 eV (i.e. 0.5 eV above ), indicating a stronger population of the conduction bands at the L- and Z-point than in the thermal model. The effect is weakened due to the overlap of the conduction band signal from the spin-orbit split core-level.
@@
This difference between the initial optical excitation and the thermalized carrier occupation can be seen in the dynamics as well: To confirm the behavior obtained from the TTM fitting, the rates from a bi-exponential rise and decay function are fitted to the transient absorption spectrum: A fast rise/decay to account for the initial thermalization of carriers, a slow decay for the carrier cooling and a damped oscillator term to account for the phonon motion (for details see Appendix H). To suppress the influence of the upper spin-orbit split core transition from the fit, a spin-orbit decomposition is applied first (see Appendix E.2). Only results where the fit successfully converges to physical measurements are reported. The resulting lifetimes are shown in Fig. 2b. On the picosecond scale, the signal relaxes with a rate (dashed line). This is due to the exponential cooling of the electron temperature that changes the energy dependent band-occupation in the Maxwell-Boltzmann limit of the Fermi-Dirac distribution (), confirming the thermal character of the signal at long timescales. Recently, Sidiropoulos et al. have connected this energy-decay scaling to a limited dimensionality of the carrier system 20. For the fast timescale, little significant spectral dependence on the thermalization rate is found. It is close to the timescale found in the component analysis (dotted line in Fig. 2b). In the valence bands, the lifetime seems to increase for energies further away from the Fermi level, although the decreased signal strength and therefore higher variance in the fit-result in this region prevents a detailed analysis. It should be noted, that the fast timescales represent a decay in the region below 31.6 eV and a rise in the region above, as evident in Fig. 2a. In the conduction band the agreement for both the slow and fast dynamics are worse, and here the fit convergence is poorer, since the transient absorption is strongly affected by the lattice dynamic. Notably, the fitted lifetimes do not correspond to previously reported lifetimes of the electronic states in Sb from DFT calculations 40, except for a small region close to the Fermi level that also corresponds to the largest reported lifetimes. This implies that the thermalization of the initially optically excited states in the experiment was not limited by the state lifetime, but appears bottlenecked, either through limited final state density at the top of the valence band or lack of open channels to dissipate excess energy through carrier-carrier or carrier-phonon scattering.
@@
In contrast to the relatively flat behavior of the thermalization timescale, a strong spectral dependence of the phase of the phonon motion is observed and shown in Fig. 3a. In the region close to the Fermi level a sharp dip in the phase is observed with the minimum occuring at the Fermi level. At lower energies, the fitted phase widely diverges due to the limited spectral range of the phonon motion. It is noted that while the phase shown here is obtained through a non-linear fit, compatible results can be obtained through Fourier analysis, as shown in Appendix J.
The phase difference of the phonon oscillation from a pure cosine form () has been associated with a mixing of the DECP mechanism with an impulsive excitation 9. In the case of Bi and Sb, O’Mahony et al. have calculated the interatomic force for different excited states 11. For the A1g phonon mode they find an initially higher force caused by the carrier population after optical excitation at 1.55 eV compared to the force after relaxation. This transient force adds an impulsive component to the displacive motion, which imparts an additional phase on the A1g mode and a reduction of the oscillation amplitude. However, such additional forces would lead to a phase of opposite sign compared to the measured phase here. It was suggested that the transient interatomic force of the optically excited carrier population could also be lower than after relaxation, e.g. in the case of As or for different photon energies in the case of Bi and Sb 11.
In the following, the approach of O’Mahony et al. is adapted to study the effect of a time-dependent phonon force that arises due to the displacive force with the rate of carrier relaxation :
where is a scaling parameter controlling the influence of carrier-relaxation. Within this model, corresponds to a limit in which only the relaxed carriers contribute to the phonon motion, i.e. at there is no force acting on the phonon. In contrast, corresponds to all carriers contributing equally, i.e. the number of excited carriers is driving the phonon motion and the constant displacive force is already established at the end of the optical excitation (see illustration in Fig. 3b).
@@
The expected phase for various values of , based on the calculated carrier relaxation rates obtained by O’Mahony et al. (see dash-dotted line in Fig. 2b) is shown in Fig 3a. Reasonable agreement with the observed phase, within the range of significantly measured phase, especially above , is seen for values of . While the minimal phase in the experiment is observed at the Fermi level, the results from DFT calculations 40 predict this minimum to occur at energies slightly below the Fermi level. Since the lifetime strongly depends on the density of states, which should be minimal at the Fermi level, this discrepancy could originate from differences in the calculated DOS.
We find the best agreement between calculated and measured phase for a value of (black dashes in Fig 3a). This indicates that indeed the relaxation of carriers is the driving force of the phonon motion. However, the measured phase clearly differs from the case that assumes only the relaxed carriers act on the phonon motion (). Indeed is not a very physical situation, as it would require the optically excited (i.e. not-relaxed) carriers to exact no force on the phonon motion. As discussed earlier, at the earliest timescales, the carrier temperature becomes ill-defined as the distribution of excited states does not follow a thermal distribution. However, as was also seen in the component analysis and TTM, a part of the carriers already occupy the same states that are occupied after thermalization, as expressed in the scaling factor ). In that sense, the obtained value of could imply the ratio of the occupation of phonon-driving states at optical excitation vs after relaxation. Following arguments 11, 31 brought forward for the driving of the Eg mode (which is not observed in our experimental configuration), this could be the difference between occupation of non fully-symmetric states (which drive the Eg mode) and their momentum-relaxation to fully-symmetric states (which would drive the A1g mode). If the momentum-relaxation drives the observed phonon phase, this would explain the differences between the spectral dependence of the observed thermalization lifetimes (Fig.2b) and the phase. A future study targeting the Eg mode could shine further light on the early-timescale forces, and investigate the differences in phase and oscillation amplitude of the A1g mode and the Eg mode, as previously suggested 11.
@@
The delay-dependent change in absorbance for the short-time-range scan is shown in Fig. 4a. As can be seen, a change in signal quickly appears around zero-delay. However, the signal does not appear to rise uniformly, but rather small, spectrally-dependent deviations from this rise are observed. Notably, in the valence band, a signal that can be associated with a strong initial hole population around the lower Z- and L-points is observed at photon energies corresponding to 1 eV to 0.5 eV below the Fermi level. The increase in absorption in this region tops out after the pump pulse is over ( 5 fs delay, see Fig. 4b), indicating direct creation of holes in this region due to resonant interband transitions. The absorption in the region closer to the Fermi level, in contrast, keeps increasing and eventually overtakes as the dominant signal, indicating a fast rearrangement of holes into this region.
@@
Interestingly, in a small band above the core-to-Fermi-level energy (32.1 eV) the signal starts to rise towards positive OD at very early delays, but continues to rise at a flatter slope compared to the signal at the lower Z-points or the Gaussian IRF (see Fig. 4c). Comparing to the signals obtained from the component analysis (cf. Fig. 1) in this region a strong contribution from the lattice dynamic is expected. It is unclear what the origin of the observed differences in rise-dynamic is. One possibility could be a different scaling between the signal that originates from carrier occupation (state blocking), which for the conduction band ( eV) would imply a trend towards negative OD, while the signal from the lattice movement through the shift of the absorption peak would trend towards positive OD. The implication would be that the start of the lattice movement does not necessarily follow the pump-pulse envelope, which would be consistent with the observation of a carrier-relaxation dependent phase of the phonon oscillation, as observed above on much longer timescales. Alternatively, a non-linear excitation process could explain the observed temporal behavior. Recent measurements on transient reflectivity in Germanium 43 has shown that different non-linear processes can show different dynamics during the envelope of the pump pulse. While no significant transitions were observed at photon energies outside of those accessible by a resonant one-photon transition, due to the narrow band gaps in the vicinity of the Fermi level, a field-induced tunneling process could for example lead to carrier excitation close to the Fermi level, as opposed to resonant one-photon interband excitations. Any non-linear excitation of carriers within the envelope of the pump pulse would not affect the measured phonon phase or decay times significantly, but could lead to an observable different dynamic at the signal onset. However, a possible explanation for the observed signals at early delays could also be a transient shifting of band-energies during the free induction decay of the core-excited states. The observed early band would imply the core-to-Fermi-level transition energy to shift towards lower energies due to a NIR induced coupling of bands. Signals from Stark-like shifting is well known from gas-phase 44, 45, 46 and core-exciton 47, 48, 49 attosecond transient absorption spectroscopy, but has so far not been observed for core-to-Fermi-level transitions in (semi-)metals.
@@
The limited signal to noise ratio in the presented data prevents a more thorough analysis, as the signal differences barely surpass the confidence interval of the measurement. An increase in sensitivity in XUV transient absorption spectroscopy, through further stabilization of the light source or referencing technique,50 would allow to further investigate the onset of the lattice dynamic on the few-femtosecond scale, e.g. by comparing the temporal dynamics seen at higher photon energies that are not overlain by the state-opening/-blocking signals or by studying the sub-cycle behavior of the photoexcitation 43.
@@ Conclusion
An XUV transient absorption spectroscopy measurement of the coherent phonon motion in antimony after displacive excitation has been shown. The results of the component analysis show the influence of the relaxation of the optically excited and thermally hot carriers on the carrier-dependent signal. The results show that the component analysis successfully separates highly transient optically excited carrier dynamics and thermalized carrier and phonon dynamics. The method obtains equivalent to complementary results as the spectral modelling fit approaches in previous studies 19, 13 with a soft-modelling approach, i.e. without an a-priori description of the underlying physical dynamics (hard-modelling). The analysis following an extended temperature model gives meaningful results, and it allows direct comparison to results obtained with other experimental techniques, even though the two temperature model breaks down at the earliest timescales, which reveals the timescale of carrier thermalization. The rich spectral information obtained by XUV transient absorption allows to obtain energy dependent relaxation, thermalization and cooling rates of carriers, as well as their effect on the lattice motion. The obtained spectral phases seem to agree well with theoretical predictions based on the relaxation of carriers, while the measured thermalization time agrees only within the limit of the theoretical predictions, implying that the thermalization is bottlenecked. Our results show how XUV transient absorption spectroscopy is able to extract meaningful timescales of electronic and structural dynamics on vastly different scales, ranging from few femtoseconds to picoseconds, from initial scattering rates of carriers, to carrier cooling times, as well as phonon decay and carrier recombination. Regarding the DECP model, the spectral dependence of the phonon oscillation phase shows that the phonon motion is not purely displacive, but is dependent on the excited state carrier lifetime to a large degree. In the language of Zeiger et al. 7, one can conclude that the DECP is driven mainly by the relaxation of carriers and therefore more akin to the change in electron temperature and not purely the number of excited carriers . This shows that carrier-relaxation needs to be considered to obtain a full picture of dynamics in solids and a pure temperature-based description is not sufficient. To optimally control the creation of large amplitude atomic motion through displacive excitation, the effects of the initial relaxation need to be considered.
@@ Appendix A Sample Growth
For the experiments, a thin film of Sb is grown on a Si3N4 membrane window (30 nm thickness, Norcada NX5050X) using a home-built physical vapor deposition oven. In a vacuum chamber, Sb pellets (99.999% purity) are mounted on a tungsten coil that is heated to about 300 °C through resistive heating. Fresh Si3N4 substrates are exposed to the Sb vapor for 6 hours. A covered part of the edge of the substrate frame during exposure allows to measure the step height of the deposited film using a commercial atomic-force scanning-tip microscope, showing a film thickness of approximately 32 nm. Sb is reported to self-crystallize to a polycrystalline form of the rhombohedral A7 phase in thin films 51 at room temperatures. A Raman spectrum of the films excited by 633 nm radiation shows the expected prominent phonon peak of the A1g mode at 148 cm-1 and Eg mode at 110 cm-1.

### v3

@@ Results
To examine the break-down of the two temperature model after the initial excitation by the pump pulse and the following carrier thermalization more closely, the measured transient absorption spectrum at early time delays are compared to the predicted spectrum from the TTM. As can be seen in Fig. 3a, the experimental data show discrepancies to the model due to the large number of non-thermalized carriers shortly after the pump pulse is over (12 fs). Compared to the model, a larger absorbance is found in a region 1 eV below the Fermi level, indicating that the pump excites carriers largely from the lower bands at the L- and Z-point (cf. Fig. 1e). On the other hand, at energies up to 0.5 eV below the Fermi level, there is less absorbance than expected from the model, indicating that fewer holes are initially in the highest valence bands. Above the Fermi level, a stronger negative change in absorbance is observed, peaking at 32.6 eV (i.e. 0.5 eV above ), indicating a stronger population of the conduction bands at the L- and Z-point than in the thermal model. The effect is weakened due to the overlap of the conduction band signal from the spin-orbit split core-level.
@@
This difference between the initial optical excitation and the thermalized carrier occupation can be seen in the dynamics as well: To confirm the behavior obtained from the TTM fitting, the rates from a bi-exponential rise and decay function are fitted to the transient absorption spectrum: A fast rise/decay to account for the initial thermalization of carriers, a slow decay for the carrier cooling and a damped oscillator term to account for the phonon motion (for details see Appendix H). To suppress the influence of the upper spin-orbit split core transition from the fit, a spin-orbit decomposition is applied first (see Appendix E.2). Only results where the fit successfully converges to physical measurements are reported. The resulting lifetimes are shown in Fig. 3b. On the picosecond scale, the signal relaxes with a rate (dashed line). This is due to the exponential cooling of the electron temperature that changes the energy dependent band-occupation in the Maxwell-Boltzmann limit of the Fermi-Dirac distribution (), confirming the thermal character of the signal at long timescales. Recently, Sidiropoulos et al. have connected this energy-decay scaling to a limited dimensionality of the carrier system [21]. For the fast timescale, little significant spectral dependence on the thermalization rate is found. It is close to the timescale found in the component analysis (dotted line in Fig. 3b). In the valence bands, the lifetime seems to increase for energies further away from the Fermi level, although the decreased signal strength and therefore higher variance in the fit-result in this region prevents a detailed analysis. It should be noted, that the fast timescales represent a decay in the region below 31.6 eV and a rise in the region above, as evident in Fig. 3a. In the conduction band the agreement for both the slow and fast dynamics are worse, and here the fit convergence is poorer, since the transient absorption is strongly affected by the lattice dynamic. Notably, the fitted lifetimes do not correspond to previously reported lifetimes of the electronic states in Sb from DFT calculations [41], except for a small region close to the Fermi level that also corresponds to the largest reported lifetimes. This implies that the thermalization of the initially optically excited states in the experiment was not limited by the state lifetime, but appears bottlenecked, either through limited final state density at the top of the valence band or lack of open channels to dissipate excess energy through carrier-carrier or carrier-phonon scattering.
@@
In contrast to the relatively flat behavior of the thermalization timescale, a strong spectral dependence of the phase of the phonon motion is observed and shown in Fig. 4a. In the region close to the Fermi level a sharp dip in the phase is observed with the minimum occurring at the Fermi level. A second minimum can be seen slightly below the Fermi level, at around -0.3 eV. These extrema coincide with extrema in the band-structure of Sb (cf. Fig. 1e). At lower energies, the fitted phase widely diverges due to the limited spectral range of the phonon motion. It is noted that while the phase shown here is obtained through a non-linear fit, compatible results can be obtained through Fourier analysis, as shown in Appendix J.
The phase difference of the phonon oscillation from a pure cosine form () has been associated with a mixing of the DECP mechanism with an impulsive excitation [9]. In the case of Bi and Sb, O’Mahony et al. have calculated the interatomic force for different excited states [11]. For the A1g phonon mode they find an initially higher force caused by the carrier population after optical excitation at 1.55 eV compared to the force after relaxation. This transient force adds an impulsive component to the displacive motion, which imparts an additional phase on the A1g mode and a reduction of the oscillation amplitude. However, such additional forces would lead to a phase of opposite sign compared to the measured phase here. It was suggested that the transient interatomic force of the optically excited carrier population could also be lower than after relaxation, e.g. in the case of As or for different photon energies in the case of Bi and Sb [11].
In the following, the approach of O’Mahony et al. is adapted to study the effect of a time-dependent phonon force that arises due to the displacive force with the rate of carrier relaxation , with a parameter that allows to test the influence of carrier-relaxation on the phonon force:
Within this model, corresponds to a limit in which only the relaxed carriers contribute to the phonon motion, i.e. at there is no force acting on the phonon. In contrast, corresponds to all carriers contributing equally, i.e. the number of excited carriers is driving the phonon motion and the constant displacive force is already established at the end of the optical excitation (see illustration in Fig. 4c).
@@
The expected phase for various values of , based on the calculated carrier relaxation rates obtained by O’Mahony et al. (see dash-dotted line in Fig. 3b) is shown in Fig 4a. Reasonable agreement with the observed phase, within the range of significantly measured phase, especially above , is seen for values of . While the minimal phase in the experiment is observed at the Fermi level, the results from DFT calculations [41] predict this minimum to occur at energies slightly below the Fermi level, but the extrema at 0 eV and -0.3 eV are otherwise well reproduced. Since the lifetime strongly depends on the density of states, which should be minimal at the Fermi level, this discrepancy could originate from differences in the calculated DOS. We find the best agreement between calculated and measured phase for a value of (black dashes in Fig 4a). This indicates that indeed the relaxation of carriers is the driving force of the phonon motion. However, the measured phase clearly differs from the case that assumes only the relaxed carriers act on the phonon motion (). Indeed is not a very physical situation, as it would require the optically excited (i.e. not-relaxed) carriers to exact no force on the phonon motion. As discussed earlier, at the earliest timescales, the carrier temperature becomes ill-defined as the distribution of excited states does not follow a thermal distribution. However, as was also seen in the component analysis and TTM, a part of the carriers already occupy the same states that are occupied after thermalization, as expressed in the scaling factor ). In that sense, the obtained value of could imply the ratio of the occupation of phonon-driving states at optical excitation vs after relaxation. Following arguments [11, 32] brought forward for the driving of the Eg mode (which is not observed in our experimental configuration), this could be the difference between occupation of non fully-symmetric states (which drive the Eg mode) and their momentum-relaxation to fully-symmetric states (which would drive the A1g mode). If the momentum-relaxation drives the observed phonon phase, this would explain the differences between the spectral dependence of the observed thermalization lifetimes (Fig.3b) and the phase. A future study targeting the Eg mode could shine further light on the early-timescale forces, and investigate the differences in phase and oscillation amplitude of the A1g mode and the Eg mode, as previously suggested [11].
@@
The delay-dependent change in absorbance for the short-time-range scan is shown in Fig. 5a. As can be seen, a change in signal quickly appears around zero-delay. However, the signal does not appear to rise uniformly, but rather small, spectrally-dependent deviations from this rise are observed. To visualize the quantitative differences of the rise signal, contour lines in Fig, 5a signify the 14%, 50% and 86% level of the maximum signal at each spectral energy, with the inset showing the corresponding delay times of the NIR pump fluence. The transient change of absorption at selected delays is shown in Fig. 5b. Notably, in the valence band, a signal that can be associated with a strong initial hole population around the lower Z- and L-points is observed at photon energies corresponding to 1 eV to 0.5 eV below the Fermi level. The increase in absorption in this region tops out after the pump pulse is over ( 5 fs delay, see Fig. 5b), indicating direct creation of holes in this region due to resonant interband transitions. The absorption in the region closer to the Fermi level, in contrast, keeps increasing and eventually overtakes as the dominant signal, indicating a fast rearrangement of holes into this region.
@@
Interestingly, in a small band above the core-to-Fermi-level energy (32.1 eV) the signal starts to rise towards positive OD at very early delays, but continues to rise at a flatter slope compared to the signal at the lower Z-points or the Gaussian IRF (see Fig. 5c). Comparing to the signals obtained from the component analysis (cf. Fig. 2) in this region a strong contribution from the lattice dynamic is expected. It is unclear what the origin of the observed differences in rise-dynamic is. One possibility could be a different scaling between the signal that originates from carrier occupation (state blocking), which for the conduction band ( eV) would imply a trend towards negative OD, while the signal from the lattice movement through the shift of the absorption peak would trend towards positive OD. The implication would be that the start of the lattice movement does not necessarily follow the pump-pulse envelope, which would be consistent with the observation of a carrier-relaxation dependent phase of the phonon oscillation, as observed above on much longer timescales. Alternatively, a non-linear excitation process could explain the observed temporal behavior. Recent measurements on transient reflectivity in Germanium [44] has shown that different non-linear processes can show different dynamics during the envelope of the pump pulse. While no significant transitions were observed at photon energies outside of those accessible by a resonant one-photon transition, due to the narrow band gaps in the vicinity of the Fermi level, a field-induced tunneling process could for example lead to carrier excitation close to the Fermi level, as opposed to resonant one-photon interband excitations. Any non-linear excitation of carriers within the envelope of the pump pulse would not affect the measured phonon phase or decay times significantly, but could lead to an observable different dynamic at the signal onset. However, a possible explanation for the observed signals at early delays could also be a transient shifting of band-energies during the free induction decay of the core-excited states, i.e. interactions with the pump-pulse during the electronic coherence of the core-excitation. The observed early band would imply the core-to-Fermi-level transition energy to shift towards lower energies due to a NIR induced coupling of bands. Signals from Stark-like shifting is well known from gas-phase [45, 46, 47] and core-exciton [48, 49, 50] attosecond transient absorption spectroscopy, but has so far not been observed for core-to-Fermi-level transitions in (semi-)metals.
Similarly, on the early timescales of the photoexcitation studied here, electronic coherences of the valence-excitation, i.e. coherence of the electron-hole pairs of the NIR-driven interband transitions, could play a role. Such coherences might appear as oscillations in delay at integer multiples of the driving field carrier period [51]. Within the results presented here we do not observe such oscillations. This could hint at a very fast dephasing of the coherences, due to carrier-carrier scattering in the presence of the high carrier density or strongly-coupled carrier-phonon scattering.
@@
The limited signal to noise ratio in the presented data prevents a more thorough analysis, as the signal differences barely surpass the confidence interval of the measurement. An increase in sensitivity in XUV transient absorption spectroscopy, through further stabilization of the light source or referencing technique,[52] would allow to further investigate the onset of the lattice dynamic on the few-femtosecond scale, e.g. by comparing the temporal dynamics seen at higher photon energies that are not overlain by the state-opening/-blocking signals or by studying the sub-cycle behavior of the photoexcitation [44].
@@ Conclusion
An XUV transient absorption spectroscopy measurement of the coherent phonon motion in antimony after displacive excitation has been shown. The results of the component analysis show the influence of the relaxation of the optically excited and thermally hot carriers on the carrier-dependent signal. The results show that the component analysis successfully separates highly transient optically excited carrier dynamics and thermalized carrier and phonon dynamics. The method obtains equivalent to complementary results as the spectral modelling fit approaches in previous studies [19, 13] with a soft-modelling approach, i.e. without an a-priori description of the underlying physical dynamics (hard-modelling). The analysis following an extended temperature model gives meaningful results, and it allows direct comparison to results obtained with other experimental techniques, even though the two temperature model breaks down at the earliest timescales, which reveals the timescale of carrier thermalization. The rich spectral information obtained by XUV transient absorption allows to obtain energy dependent relaxation, thermalization and cooling rates of carriers, as well as their effect on the lattice motion. The obtained spectral phases seem to agree well with theoretical predictions based on the relaxation of carriers, while the measured thermalization time agrees only within the limit of the theoretical predictions, implying that the thermalization is bottlenecked. Our results show how XUV transient absorption spectroscopy is able to extract meaningful timescales of electronic and structural dynamics on vastly different scales, ranging from few femtoseconds to picoseconds, from initial scattering rates of carriers, to carrier cooling times, as well as phonon decay and carrier recombination. Regarding the DECP model, the spectral dependence of the phonon oscillation phase shows that the phonon motion is not purely displacive, but is dependent on the excited state carrier lifetime to a large degree. In the language of Zeiger et al. [7], one can conclude that the DECP is driven mainly by the relaxation of carriers and therefore more akin to the change in electron temperature and not purely the number of excited carriers . This shows that carrier-relaxation needs to be considered to obtain a full picture of dynamics in solids and a pure temperature-based description is not sufficient. To optimally control the creation of large amplitude atomic motion through displacive excitation, the effects of the initial relaxation need to be considered. To further understand the involved dynamics, powerful theoretical methods that couple the ultrafast carrier dynamics as well as the phonon motion are required. Interesting further experimental studies could investigate the efficiency of quenching of the phonon motion by a second NIR pulse, similar as recently done for charge density wave (CDW) materials [53] and studying the effect of spectral phase shaping of the pump pulses on the phonon phases. The role of carrier relaxation is thus expected to be important for many phenomena that underlie manipulation of Peierls distorted materials and optically-driven phase transitions via photoexcitation of carriers, including for example switching of charge-density waves.
@@ Acknowledgments
Data Availability
Experimental and computational data that support the findings in this study are available on Zenodo [54].
@@ Appendix A Sample Growth
For the experiments, a thin film of Sb is grown on a Si3N4 membrane window (30 nm thickness, Norcada NX5050X) using a home-built physical vapor deposition oven. In a vacuum chamber, Sb pellets (99.999% purity) are mounted on a tungsten coil that is heated to about 300 °C through resistive heating. Fresh Si3N4 substrates are exposed to the Sb vapor for 6 hours. A covered part of the edge of the substrate frame during exposure allows to measure the step height of the deposited film using a commercial atomic-force scanning-tip microscope, showing a film thickness of approximately 32 nm. Sb is reported to self-crystallize to a polycrystalline form of the rhombohedral A7 phase in thin films [55] at room temperatures. A Raman spectrum of the films excited by 633 nm radiation shows the expected prominent phonon peak of the A1g mode at 148 cm-1 and Eg mode at 110 cm-1.